OpenStudy (anonymous):

Greg has only nickels and dimes in his money box. He knows that he has more than $10 in the box. Let x represent the number of nickels in the box and y represent the number of dimes in the box. Which of the following statements best describes the steps to graph the solution to the inequality in x and y?

OpenStudy (anonymous):

Which of the following statements best describes the steps to graph the solution to the inequality in x and y? Draw a dashed line to represent the graph of 5x + 10y = 1000, and shade the portion below the line for positive values of x and y. Draw a dashed line to represent the graph of 5x + 10y = 1000, and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 10x + 5y = 1000, and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 10x + 5y = 1000, and shade the portion below the line for positive values of x and y.

OpenStudy (anonymous):

@iambatman , @KendrickLamar2014 , @just_one_last_goodbye , @sweetburger , @sissyedgar , @SyedMohammed98 , @Rainbow_dashie , @robtobey , @saifoo.khan , @abb0t , @goformit100

OpenStudy (anonymous):

@bohotness , @iGreen